OpenStudy (anonymous):

(sq. root sign) x+3=10

hartnn (hartnn):

is it \(\sqrt x+3=10 \\or,\sqrt {x+3}=10 \)

OpenStudy (anonymous):

the second one

hartnn (hartnn):

square both sides, what u get ?

OpenStudy (anonymous):

so it would be 103?

hartnn (hartnn):

did you get x+3 =100 ?

OpenStudy (anonymous):

yes and the add three to both sides and x = 103? right

hartnn (hartnn):

no, x+3 =100 to isolate 'x' , we need to SUBTRACT 3 from both sides, then what u get ?

OpenStudy (anonymous):

oh okay so it is 97

OpenStudy (anonymous):

that is where I get confused is whether to add or subtract

hartnn (hartnn):

yes.

hartnn (hartnn):

x+3 =100 since 3 is ADDED to x, we do inverse operation of SUBTRACTION to isolate x since, x+3-3 = x+0 = x

OpenStudy (anonymous):

oh okay thank you I think I got it

hartnn (hartnn):

welcome ^_^